# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v2

@@ 3 Literature Review Methodology
To provide a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [41, 42]. While this is not a formal systematic review, we followed a transparent and reproducible process to ensure coverage and rigor. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
@@
3.1 Research Objectives
The review was guided by the following four specific questions:
@@ 3.1 Research Objectives
3.2 Data Sources and Search Strategy
To conduct a thorough literature review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
(“ Agentic AI” OR “multi-agent systems” OR “MAS” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
@@ 3.2 Data Sources and Search Strategy
This comprehensive search strategy ensured the identification of relevant studies addressing the integration of TRiSM principles with LLM-based multi-agent systems, particularly in high-stakes domains such as healthcare, science, and finance.
3.3 Inclusion and Exclusion Criteria
We applied the following criteria during the screening process:
  • Inclusion:
@@ 3.3 Inclusion and Exclusion Criteria
3.4 Quality Assurance
To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [43, 42]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
@@ 4 Fundamentals of Agentic AI Systems
4.1 Agentic AI vs. Traditional AI Agents
Agentic AI systems fundamentally transform the design of autonomous systems, and are different from traditional AI agents. Traditional agents operate through predefined rules [44], heuristic workflows [45, 46], or deterministic logic [47], and perform well in narrow, well-defined environments. Cognitive Agents [48], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [49] or ACT-R [50]. However, they remain constrained by static knowledge bases and limited adaptability [48].
@@ 5 The TRiSM Framework for Agentic AI
The TRiSM Framework for Agentic AI is presented in this section.
@@
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) [89]
@@
5.1 Governance for Agentic LLM Systems
AI TRiSM is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards [29]. Originally highlighted in industry guidelines for AI governance [93], TRiSM provides a structured approach to manage the unique challenges of LLM-based “ Agentic AI" systems (herein we refer to as Agentic LLM systems). The TRiSM framework addresses the trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 29]. Figure 4 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
@@ 5.1 Governance for Agentic LLM Systems
The TRiSM framework aligns closely with several key international AI governance efforts that shape best practices for safety, accountability, and transparency in AI systems. For example, the EU AI Act [94] (notably Article 14) mandates clear obligations for transparency, traceability, and human oversight in high-risk systems, principles directly reflected in TRiSM Explainability and ModelOps pillars. The NIST AI Risk Management Framework (AI RMF) [95] emphasizes Explainability, Governability, and Validity as core risk mitigation strategies, providing a conceptual match to TRiSM’s structured controls across lifecycle and system interaction points. Additionally, the OECD AI Principles advocate for transparency, robustness, and accountability [96]each echoed in TRiSM attention to secure tool invocation, memory traceability, and multi-agent auditability. These global frameworks reinforce the need for Agentic AI systems to support real-time monitoring, enforce role-based access, maintain verifiable logs, and provide intelligible outputs. As shown in Figure 5, TRiSM pillars provide operational scaffolding for implementing regulatory expectations in LLM-based multi-agent contexts.
@@
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [94] and the NIST AI RMF “Explain” function [95]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [97]. ModelOps (Lifecycle Governance): Anchored in the NIST RMF “Govern” function [95] and ISO/IEC 42001 [98]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [90] and OECD AI Principles on Robustness [96], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [91], HIPAA §164 [92], and OECD AI Principles [96], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
@@
Figure 5: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
5.2 Explainability in Multi-Agent Decision Making
@@
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of AMAS, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [99]. However, achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
@@
Established explainable AI techniques provide a starting point. For instance, LIME [100] and Shapley Additive Explanations (SHAP) [101] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
@@
Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered [102], for example, if a particular agent’s contribution were removed or modified. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making. Recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g. CoT prompts or dialogue between agents) can further enhance transparency. Human auditors can also help with a trace of how agents arrived at a decision, e.g., which agent contributed what information and why, to produce narrative explanations for its outcomes.
5.3 ModelOps : Lifecycle Management for LLM Agents
ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [103]. It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production. Within Agentic AI systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps is important to maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
@@ 5.3 ModelOps : Lifecycle Management for LLM Agents
Agentic AI systems require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions. This aligns with best practices in LLM operations (LLMOps) [104], which integrate MLOps principles tailored to the challenges of LLMs. In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
5.4 Application Security for LLM Agents
@@
5.5 Model Privacy and Data Protection
The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [27]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. Therefore, TRiSM’s privacy pillar recommends organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
@@ 5.5 Model Privacy and Data Protection
To preserve privacy in multi-agent LLM systems, several technical measures are employed. Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [109]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [110]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [111]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [112]. Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [113]. Finally, the implementation of model privacy policies and compliance frameworks—including data retention limits, audit logs, and adherence to regulations such as GDPR or HIPAA, ensures that agents operate within legal and ethical boundaries [114].
Table 4 condenses each pillar into its governing goal.
6 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
In this section, we present the main threats in Agentic AI and also discuss a taxonomy of emergent risk classes.
6.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [115]. Below, we discuss these threats.
Autonomy abuse. The foremost threat is autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs [116]. Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states [117].
Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential adversarial injections [118, 119]. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms [116].
Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [19] and AutoGen [18]. These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
6.2 Taxonomy of Risks
We propose a taxonomy of risks associated with autonomous multi-agent systems (AMAS), categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 6. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [120, 121]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [17]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information [122, 123]. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [124, 125]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [126, 127]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [128, 129]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [130], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals [131]. These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
6.3 Real-World Use Cases
Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental Agentic AI systems.
Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [132], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [133]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks [134].
7 Trust and Explainability in Agentic AI

### v3

@@ 3 Literature Review Methodology
To review the literature related to TRiSM as applied in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [39, 40]. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
@@
Research Objectives The review was guided by the following four specific questions:
@@
Search Strategy To conduct a review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
(“ Agentic AI” OR “multi-agent systems” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
@@
Inclusion and Exclusion Criteria We applied the following criteria during the screening process:
Inclusion Criteria:
@@
Quality Assurance To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [41]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
@@ 4 Fundamentals of Agentic AI Systems
4.1 Traditional AI Agents vs Agentic AI
Traditional agents operate through predefined rules [42], heuristic workflows [43, 44], or deterministic logic [45], and perform well in narrow, well-defined environments. Cognitive Agents [46], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [47] or ACT-R [48]. However, they remain constrained by static knowledge bases and limited adaptability [46].
@@ 4.2 LLM-Based Agentic AI System Architecture
5 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
5.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [87]. Below, we discuss some of these major threats.
  • Autonomy abuse. The foremost threat is the autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs . Unlike deterministic models, these Agentic AI systems dynamically generate actions, complicating efforts to define and enforce safe operational states .
  • Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential prompt injections. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms .
  • Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT and AutoGen . These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
  • Goal Misalignment and Exposure. Agentic AI systems often operate with internalized or user-specified goals. Improper scoping or exposure of goal representations can lead to prompt extraction, misalignment, or competitive agent behaviors, particularly when third-party tools or APIs are invoked dynamically during task completion.
  • Tool Misuse and External API Exploits. Agents capable of invoking external tools (e.g., via APIs, code execution, or web access) present novel attack surfaces. Misuse of these tools, intentional or accidental, can trigger financial costs, Distributed Denial of Service (DDoS) behavior, or legal violations. Sandboxing and invocation tracing remain limited in most open-source frameworks.
  • Multi-Agent Collusion or Drift. Emergent coordination among agents, while useful for collaboration, can also lead to collusion, competitive drift, or polarization of outputs, especially in self-reinforcing memory loops or agent swarms. Such behavior complicates interpretability and may undermine safety alignment goals.
5.2 Taxonomy of Risks
We propose a taxonomy of risks associated with AMAS, categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 4. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
  • Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [88]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [13]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
  • Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [89]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
  • Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [90]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
  • Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [91]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [92], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals . These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
5.3 Real-World Examples
Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [93], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [94]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks.
Next, we discuss the AI TRiSM prism framework and map its core principles to a AMAS.
@@ 6 The TRiSM Framework for Agentic AI
6.1 Overview of AI TRiSM and Its Core Components
The TRiSM Framework, short for Trust, Risk, and Security Management, offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
@@ 6.1 Overview of AI TRiSM and Its Core Components
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
@@
  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness. Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [50]. This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies. This includes technical risks such as algorithmic bias [37], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [32]. Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access. This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [19]. Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
Originally highlighted in industry guidelines for AI governance [105], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems. The TRiSM framework addresses trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 23]. Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
@@
6.2 Mapping Pillars of TRiSM with Agentic AI
@@
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [106] and the NIST AI RMF “Explain” function [107]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [108]. ModelOps : Anchored in the NIST RMF “Govern” function [107] and ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [102] and OECD AI Principles on Robustness [109], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [103], HIPAA §164 [104], and OECD AI Principles [109], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds. Governance: Grounded in overarching regulatory frameworks such as GDPR, CCPA [110], and the OECD AI Principles [109], this pillar supports institutional oversight, policy compliance, and cross-system accountability. It enforces auditable workflows, verifiable role-based access, regulatory traceability, and human oversight, particularly in multi-agent and high-risk AI deployments.
@@
Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, 42005, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
Explainability in Multi-Agent Decision Making
@@
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of LLM-based multi-agent systems, it is paramount for building user trust because outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [111]. However, achieving explainability in such systems is particularly challenging , one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to the final outcomes [112]. In other words, explaining why a multi-agent system behaved a certain way requires tracing which agent did what and how their collaboration produced an emergent result, a task that traditional single-model explainability techniques do not directly address. Several approaches are being explored to enhance explainability in LLM-driven multi-agent systems:
@@
  • Feature Attribution (LIME/ SHAP): Established explainable AI techniques provide a starting point. For instance, LIME) [113] and Shapley Additive Explanations (SHAP) [114] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent. By highlighting the most influential parts of each agent’s input, methods like LIME/SHAP help clarify individual agent decisions in the multi-agent mix [115].
@@
  • Counterfactual Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [116]. For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency. Logging the step-by-step reasoning process – for example, an agent’s chain-of-thought (CoT) prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [69]. By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision. This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
  • Natural Language Explanations: Finally, the generative nature of LLMs can be leveraged to turn the above insights and logs into human-friendly explanations [117]. Instead of expecting end-users to read raw CoT logs or analyze feature weights, an “explainer” agent (which could be another LLM) or a human analyst can compile the reasoning trace into a narrative. In fact, an important aspect of explainability is communicating the reasons behind decisions in an intelligible way. By using natural language, the system can answer user questions like “Why did the agents decide to do X?” in a conversational manner.
ModelOps : Lifecycle Management for LLM Agents
@@
ModelOps (Machine Learning Operations) is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [118]. Within AMAS, ModelOps covers not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps helps maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system. AMAS require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
Application Security for LLM Agents
@@
Model Privacy and Data Protection
@@
The model privacy addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [21]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. To preserve privacy in AMAS, several technical measures can be employed. For example, Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [123]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [124]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [125]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [126].Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [127].
Governance
Governance of AI systems necessitates compliance with evolving legal and policy frameworks, including the General Data Protection Regulation (GDPR) [103] and the California Consumer Privacy Act (CCPA) [110]. Effective governance extends beyond regulatory adherence to cover proactive management of legal, ethical, and societal risks. This includes establishing transparent accountability structures, maintaining auditable documentation, and ensuring the protection of data subject rights throughout the AI lifecycle.
Several global frameworks provide guiding principles that align with the TRiSM pillars. The EU AI Act [106], particularly Article 14, mandates transparency, traceability, and human oversight in high-risk systems; principles embedded within TRiSM’s Explainability and ModelOps dimensions. Similarly, the NIST AI Risk Management Framework (AI RMF) [107] identifies: Explainability, Governability, and Validity as key components of AI risk mitigation, which directly correspond to TRiSM’s lifecycle control mechanisms. The OECD AI Principles [109] reinforce the importance of robustness, transparency, and accountability, echoing TRiSM’s attention to secure tool invocation, memory traceability, and multi-agent auditability.
Together, these frameworks underscore the need for governance architectures in Agentic AI systems that support real-time monitoring, role-based access, verifiable logs, and intelligible outputs. As illustrated in Figure 6, TRiSM offers a practical scaffolding for operationalizing these requirements in LLM-based multi-agent environments.
7 Trust and Explainability in Agentic AI Systems